OpenStudy (anonymous):

Which of the following is a solution to the equation log 4 x + log 4 (x – 3) = 1 ?

sam (.sam.):

\[\log _4(x)+\log _4(x-3)=1\]

sam (.sam.):

Using rule \(\large \log_a b=\frac{\log_cb}{\log_ca}\) \[\frac{\log (x-3)}{\log (4)}+\frac{\log (x)}{\log (4)}=1\] Combine \[\frac{\log ((x-3) x)}{\log (4)}=1\] Multiply both sides by log(4): \[\log((x-3) x)=\log(4)\] \[(x-3) x=4\] \[x^2-3 x=4\] \[x^2-3x-4=0\] Solve